The following is a set of operational information about a local motel during the past years.   The motel has 200 rooms in its property. It operates 365 days every year.  (Ignore the leap year in the calculation process.)  Please be advised that the Year "n" is the most recent year.

Year          Occupancy Rate                  ADR                      

n - 5              70%                                    $120                      

n - 4              72%                                  $130      

n - 3              69%                                    $170                      

n - 2              70%                                    $168

n - 1              68%                                    $174      

n             72%                                  $180

n + 1

n + 2

1. Using the method of 3-Year Moving Average, project the annual room revenues ($) of the motel above for the next year (Years "n + 1"). The number of days of operations and the number of rooms available for sales will remain the same (200) for the next few years. Again, ignore the leap year. All years have 365 operating days.

2. Based on your findings on the previous question; using the method of 3-Year Moving Average, project the annual rooms revenues ($) of the motel above for the second year ("n + 2") from the current year ("n"). The number of days of operations and the number of rooms available for sales will remain the same (200) for the next few years. Again, ignore the leap year. All years have 365 operating days.
Answer #1

Answer :

Calculation of moving average of occupancy rate and rate

3 - years moving average of occupancy = 72 + 68 + 70 / 3 = 70 %

70% - $174

100% - ?

For 100% of occupancy = 174 *100/70 = 248.57 or $249

Hence on n+1 year for 70% of occupancy estimated revenue of $174

Calculation of n+2

n-1 68% $174

n 72% $180

n+1 70% $174

Occupancy in n+2 year is = 70 +72 +68 /3 =70%

Revenue in n+2 year is = 174 + 180 + 174 /3 = $176

On n+2 year for 70% of occupancy estimated revenue us $176.

Please Upvote Thank You!!!